5x+2x-9=40 we just need to solve the problem.
Answer by jai_kos(139)

Given 5x + 2x - 9 = 40
Group up all the "x" terms in one side and all the numbers on the other side,
We get
5x +2x = 40 + 9
7x = 49
Divide by 7 on both sides, we get
x = 49 / 7
x = 7